Here's the question... A rectangular prism has a square base with an area 4 mē. The surface area of the prism is 48 mē. What are the dimensions of the prism?
Answer by rothauserc(4718)

surface area of a rectangular prism =
2*( wh + lw + lh) where l is length, w is width, and h is height
note that l = w = 2 since the base is a square whose area is 4
therefore
48 = 2*( 2h + 4 + 2h)
48 = 4h + 8 + 4h
8h = 40
h = 5
l = 2m, w = 2m, h = 5m
**********************************************
check answer
48 = 2 * ( (2*5) + (2*2) + 2*5) )
48 = 2 * ( 10 + 4 + 10 )
48 = 2 * 24
48 = 48
our answer checks